# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ 5 Experiment
In this section, we evaluate the robustness of proposed PRISM framework to noisy demonstrations in long-horizon environments and analyze how skill discovery and refinement during training enhance performance.
@@ 5.1 Experimental Setup
We compare the proposed PRISM with 3 non-meta RL baselines: SAC, which trains test tasks directly without using the offline dataset; SAC+RND, which incorporates RND-based intrinsic noise for enhanced exploration; and SPiRL, which learns skills from the offline dataset using Eq. (1) and trains high-level policies for individual tasks. Also, we include 4 meta-RL baselines: PEARL, a widely used context-based meta-RL algorithm without skill learning; PEARL+RND, which integrates RND-based exploration into PEARL; SiMPL, which applies skill-based meta-RL using Eq. (2); and our PRISM. PRISM’s hyperparameters primarily follow Nam et al. (2022), with additional parameters (e.g., temperature $T$) tuned via hyperparameter search, while other baselines use author-provided code. Results are averaged over 5 random seeds, with standard deviations represented as shaded areas in graphs and $\pm$ values in tables.
@@ 5.2 Environmental Setup
Environment(Noise) SAC SAC+RND PEARL PEARL+RND SPiRL SiMPL PRISM
Kitchen(Expert) $0.01$$\pm$$0.01$ $0.02$$\pm$$0.05$ $0.23$$\pm$$0.14$ $0.42$$\pm$$0.16$ $3.11$$\pm$$0.33$ $3.40$$\pm$$0.18$ $\mathbf{3.97}$$\pm$$0.09$
Kitchen($\sigma=0.1$) $3.37$$\pm$$0.31$ $3.76$$\pm$$0.14$ $\mathbf{3.91}$$\pm$$0.12$
Kitchen($\sigma=0.2$) $2.06$$\pm$$0.43$ $2.18$$\pm$$0.33$ $\mathbf{3.73}$$\pm$$0.16$
Kitchen($\sigma=0.3$) $0.83$$\pm$$0.17$ $0.81$$\pm$$0.25$ $\mathbf{3.48}$$\pm$$0.07$
Office(Expert) $0.00$$\pm$$0.00$ $0.00$$\pm$$0.00$ $0.01$$\pm$$0.01$ $0.01$$\pm$$0.01$ $0.65$$\pm$$0.24$ $2.50$$\pm$$0.26$ $\mathbf{2.86}$$\pm$$0.35$
Office($\sigma=0.1$) $0.91$$\pm$$0.31$ $3.33$$\pm$$0.39$ $\mathbf{3.40}$$\pm$$0.38$
Office($\sigma=0.2$) $0.49$$\pm$$0.22$ $1.20$$\pm$$0.24$ $\mathbf{2.01}$$\pm$$0.24$
Office($\sigma=0.3$) $0.42$$\pm$$0.14$ $0.11$$\pm$$0.04$ $\mathbf{1.68}$$\pm$$0.15$
Maze2D(Expert) $0.20$$\pm$$0.06$ $0.35$$\pm$$0.07$ $0.10$$\pm$$0.01$ $0.11$$\pm$$0.08$ $0.77$$\pm$$0.06$ $0.80$$\pm$$0.04$ $\mathbf{0.87}$$\pm$$0.05$
Maze2D($\sigma=0.5$) $\mathbf{0.89}$$\pm$$0.03$ $0.87$$\pm$$0.05$ $\mathbf{0.89}$$\pm$$0.03$
Maze2D($\sigma=1.0$) $0.80$$\pm$$0.01$ $0.87$$\pm$$0.05$ $\mathbf{0.93}$$\pm$$0.05$
Maze2D($\sigma=1.5$) $0.81$$\pm$$0.05$ $0.68$$\pm$$0.06$ $\mathbf{0.99}$$\pm$$0.02$
AntMaze(Expert) $0.00$$\pm$$0.00$ $0.00$$\pm$$0.00$ $0.00$$\pm$$0.00$ $0.00$$\pm$$0.00$ $0.64$$\pm$$0.09$ $0.67$$\pm$$0.07$ $\mathbf{0.81}$$\pm$$0.08$
AntMaze($\sigma=0.5$) $0.76$$\pm$$0.10$ $0.77$$\pm$$0.05$ $\mathbf{0.82}$$\pm$$0.05$
AntMaze($\sigma=1.0$) $0.50$$\pm$$0.06$ $0.33$$\pm$$0.09$ $\mathbf{0.60}$$\pm$$0.02$
AntMaze($\sigma=1.5$) $0.30$$\pm$$0.01$ $0.27$$\pm$$0.05$ $\mathbf{0.41}$$\pm$$0.01$
We evaluate algorithms across 4 long-horizon, multi-task environments: Kitchen and Maze2D from Nam et al. (2022), and Office and AntMaze, newly introduced in this work, as illustrated in Fig. 5. Offline datasets $\mathcal{B}_{\text{off}}$ are generated by policies perturbed from the expert policy using Gaussian action noise at varying levels $\sigma$, tailored to each environment. Detailed descriptions are as follows:
Kitchen: The Franka Kitchen environment in D4RL benchmark (Fu et al., 2020b), proposed by Gupta et al. (2020), features a robotic arm completing 4 sequential subtasks selected from 7 available subtasks, receiving a reward of 1 for each successful subtask. The environment includes 25 meta-train and 10 meta-test tasks, with noise levels ranging from expert demonstrations (noise-free) to Gaussian action noise at $\sigma=0.1$, $0.2$, and $0.3$.
Office: The Office environment, adapted from Pertsch et al. (2022) for meta-RL. Each task requires moving 3 randomly selected office objects into one of 3 containers on a desk, receiving a reward of 1 for each successful pick or place action. The environment includes 7 objects, 25 meta-train tasks, and 10 meta-test tasks, with noise levels ranging from expert to Gaussian noise with $\sigma=0.1$, $0.2$, and $0.3$.
Maze2D: Based on D4RL (Fu et al., 2020b), this navigation task requires a 2-DoF point mass agent to reach a goal point in a larger 20x20 maze than the default D4RL Maze2D, receiving a reward of 1 upon reaching the goal. The environment includes 40 meta-train tasks and 10 meta-test tasks, with noise levels ranging from expert to Gaussian action noise at $\sigma=0.5$, $1.0$, and $1.5$.
AntMaze: Similar to Maze2D from D4RL (Fu et al., 2020b), this environment uses a more complex ant agent in a smaller 10x10 maze to account for the increased difficulty, receiving a reward of 1 upong reaching the goal. The environment includes 20 meta-train tasks and 10 meta-test tasks, with noise levels ranging from expert to Gaussian action noise at $\sigma=0.5$, $1.0$, and $1.5$.
More detailed experimental setup, including explanations of other baselines, descriptions of the considered environments and data collection processes, as well as our hyperparameter configurations, can be found in Appendix B.
@@ 5.3 Performance Comparison
We compare the proposed PRISM with various baseline algorithms. Non-meta RL algorithms are trained directly on each test task for 0.5K iterations due to the absence of a meta-train phase. Meta-RL algorithms undergo meta-training for 10K iterations in Kitchen and Office, and 4K in Maze2D and AntMaze, followed by fine-tuning on test tasks for an additional 0.5K iterations. For PRISM, the skill update interval $K_{\text{iter}}$ is set to 2K for Kitchen, Office, and AntMaze; 1K for Maze2D. To ensure a fair comparison, PRISM counts each update process from its exploration and high-level policies as one iteration. Table 1 presents the final average return across test tasks after the test iterations. From the result, SAC and PEARL baselines, which do not utilize skills or offline datasets, perform poorly on long-horizon tasks, yielding a single result across all noise levels. In contrast, SPiRL, SiMPL, and PRISM, which leverage skills, achieve significantly better performance.
Figure 6: Learning curves of the meta-train and meta-test phases on Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). PRISM ($\pi_{\text{exp}}$) and PRISM ($\pi_{h}$) denote the performance of the exploration policy $\pi_{\text{exp}}$ and high-level policy $\pi_{h}$ during meta-training.
@@
SPiRL and SiMPL, however, show sharp performance declines as dataset noise increases. While both perform well with expert data, SiMPL struggles under noisy conditions due to instability in its task encoder $q_{e}$, sometimes performing worse than SPiRL. Here, the baseline results for Maze2D (Expert) are somewhat lower than those reported in the SiMPL paper. This discrepancy likely arises because, in constructing the offline dataset, we considered fewer tasks compared to SiMPL, resulting in trajectories that do not fully cover the map. Interestingly, minor noise occasionally boosts performance by introducing diverse trajectories that improve skill learning. In contrast, PRISM demonstrates superior robustness across all evaluated environments, consistently outperforming baselines at varying noise levels. For example, in the Kitchen environment, PRISM maintains strong performance under significant noise by effectively refining useful skills, while in Maze2D, higher noise levels lead to the discovery of diverse skills, achieving perfect task completion when $\sigma=1.5$. These results highlight PRISM’s ability to refine and discover robust skills, significantly enhancing meta-RL performance. Moreover, PRISM excels with both noisy and expert data, achieving superior test performance by learning more effective skills.
Fig. 6 shows the learning progress during meta-train and meta-test phases for Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), emphasizing the significant performance gap between PRISM and other methods. The periodic drop in PRISM’s high-level performance corresponds to the reinitialization of $\pi_{h}$ every $K_{\text{iter}}$. Non-meta RL algorithms, even those with RND-based exploration, struggle with the considered long-horizon tasks. Similarly, SPiRL and SiMPL, constrained by noisy offline datasets, exhibit limited improvement due to reliance on fixed offline data. In contrast, PRISM periodically refines skills through prioritized skill refinement, leading to superior meta-test performance. Additional results for other noise levels and the evolution of control factor $\beta$ are provided in Appendix C.
@@
Figure 7: Visualization of buffer control factor $\beta$ dynamics and refined skill evolution in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). In Kitchen, the refined skills at $t=15$ and $t=30$ during the microwave-opening task are depicted, while in Maze2D, trajectories using refined skills illustrate the process of progressively expanding to broader areas to solve tasks.
5.4 Visualization of Skill Refinement Process
Figure 8: Ablation study: Component evaluation
Figure 9: Ablation study: Prioritization temperature $T$
To analyze skill learning and refinement, Fig. 7 shows the evolution of the buffer control factor $\beta$ and skill refinement in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). For Kitchen, the microwave-opening subtask is examined, while for Maze2D, skill improvements are observed in navigating the maze. In the early stages (1K iterations for Kitchen, 0.5K for Maze2D), pretrained skills learned solely from the offline dataset, are used as no skill updates occur until $K_{\text{iter}}$. This results in poor performance: in Kitchen, the agent fails to grasp the microwave handle, and in Maze2D, noisy trajectories lead to task failures. As training progresses, $\beta$ initially relies on offline data but gradually increases as high-quality samples accumulate in the online buffer. This shift enables task-relevant skill refinement and improved task-solving.
@@ 5.4 Visualization of Skill Refinement Process
As a result, by iteration 5K in Kitchen, the agent learns to open the microwave, refining this skill to complete the task more efficiently by iteration 10K. In Maze2D, the agent explores more diverse trajectories over iterations, ultimately solving all training tasks by iteration 4K. These results highlight how PRISM refines skills iteratively by leveraging prioritized data from offline and online buffers. As shown in Table 1, the learned skills generalize effectively to unseen test tasks, demonstrating PRISM’s robustness and efficacy. In addition, improvement in task representation through skill refinement, along with additional visualizations, including skill representations, are provided in Appendix D.
@@ 5.5 Ablation Studies
We evaluate the impact of PRISM’s components and key hyperparameters in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), focusing on the effect of the prioritization temperature $T$. Additional analyses, including component evaluation and KLD coefficient $\lambda^{\text{kld}}_{\text{exp}}$ for all noise levels, are in Appendix E.
Component Evaluation: To evaluate the importance of PRISM’s components, we compare the meta-test performance of PRISM with all components included against the following variations: (1) Without $\mathcal{B}_{\text{off}}$, relying solely on $\mathcal{B}_{\text{on}}^{i}$; (2) Without $P_{\mathcal{B}_{\text{off}}}$, applying uniform sampling in $\mathcal{B}_{\text{off}}$ instead of maximum return relabeling; (3) Without $\mathcal{B}_{\text{on}}$, using only $\mathcal{B}_{\text{off}}$; and (4) Without $\pi_{\text{exp}}$, removing the exploration policy. As illustrated in Fig. 9, performance drops significantly when either buffer is removed, showing that both are crucial for effective skill discovery. Uniform sampling in $\mathcal{B}_{\text{off}}$ also reduces performance, underlining the importance of maximum return relabeling. Lastly, excluding $\pi_{\text{exp}}$ notably degrades results, emphasizing the critical role of exploration in skill refinement and trajectory discovery.
Prioritization Temperature $T$: The prioritization temperature $T$ adjusts the prioritization between online and offline buffers. Specifically, lower $T$ biases sampling toward high-return buffers, while higher $T$ results in uniform sampling. As shown in Fig. 9, meta-test performance varies with $T$. When $T=0.1$, performance drops due to an excessive focus on one buffer, as the trends observed in the component evaluation. Conversely, high $T=2.0$ also degrades performance by eliminating prioritization. These results highlight the importance of proper tuning: $T=1.0$ for Kitchen and $T=0.5$ for Maze2D achieve the best performance.

### v3

@@ 5 Experiment
In this section, we evaluate the robustness of the SISL framework to noisy demonstrations in long-horizon environments and analyze how self-improving skill learning enhance performance.
@@ 5.1 Experimental Setup
We compare the proposed SISL with 3 non-meta RL baselines: SAC, which trains test tasks directly without using the offline dataset; SAC+RND, which incorporates RND-based intrinsic noise for enhanced exploration; and SPiRL, which learns skills from the offline dataset using Eq. (1) and trains high-level policies for individual tasks. Also, we include 4 meta-RL baselines: PEARL, a widely used context-based meta-RL algorithm without skill learning; PEARL+RND, which integrates RND-based exploration into PEARL; SiMPL, which applies skill-based meta-RL using Eq. (2); and our SISL. SISL’s hyperparameters primarily follow Nam et al. (2022), with additional parameters (e.g., temperature $T$) tuned via hyperparameter search, while other baselines use author-provided code. Results are averaged over 5 random seeds, with standard deviations represented as shaded areas in graphs and $\pm$ values in tables.
@@ 5.2 Environmental Setup
Environment(Noise) SAC SAC+RND PEARL PEARL+RND SPiRL SiMPL SISL
[…]
Maze2D(Expert) $20±\pm0.0$60.06 $35±\pm0.0$70.07 $\pm0.010.01$ $11±\pm0.0$80.08 $77±\pm0.0$60.06 $80±\pm0.0$40.04 $\mathbf{0.87}$$\pm$$0.05$
Maze2D($\sigma=0.5$) $\mathbf{0.89}$$\pm$$0.03$ $87±\pm0.0$50.05 $\mathbf{0.89}$$\pm$$0.03$
Maze2D($\sigma=1.0$) $80±\pm0.0$10.01 $87±\pm0.0$50.05 $\mathbf{0.93}$$\pm$$0.05$
Maze2D($\sigma=1.5$) $81±\pm0.0$50.05 $68±\pm0.0$60.06 $\mathbf{0.99}$$\pm$$0.02$
AntMaze(Expert) $\pm0.000.00$ $\pm0.000.00$ $\pm0.000.00$ $\pm0.000.00$ $64±\pm0.0$90.09 $67±\pm0.0$70.07 $\mathbf{0.81}$$\pm$$0.08$
[…]
AntMaze($\sigma=1.0$) $50±\pm0.0$60.06 $33±\pm0.0$90.09 $\mathbf{0.60}$$\pm$$0.02$
AntMaze($\sigma=1.5$) $30±\pm0.0$10.01 $27±\pm0.0$50.05 $\mathbf{0.41}$$\pm$$0.01$
Figure 6: Learning curves of the meta-train and meta-test phases on Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). SISL ($\pi_{\text{imp}}$) and SISL ($\pi_{h}$) denote the performance of the skill-improvement policy $\pi_{\text{imp}}$ and high-level policy $\pi_{h}$ during meta-training.
We evaluate SISL across four long-horizon, multi-task environments: Kitchen and Maze2D from Nam et al. (2022), and Office and AntMaze, newly introduced in this work, as illustrated in Fig. 5. Offline datasets $\mathcal{B}_{\text{off}}$ are generated by perturbing expert policies with varying levels of Gaussian action noise, tailored to each environment. In the Kitchen environment, based on the Franka Kitchen from the D4RL benchmark (Fu et al., 2020b) and proposed by Gupta et al. (2020), a robotic arm completes a sequence of subtasks, with noise levels ranging from expert to $\sigma=0.1$, $0.2$, and $0.3$. The Office environment, adapted from Pertsch et al. (2022), involves picking and placing randomly selected objects into containers, with Gaussian noise applied at the same levels as Kitchen. Maze2D, based on D4RL (Fu et al., 2020b), requires a point-mass agent to navigate a large 20x20 maze, while AntMaze features a more complex ant agent maneuvering through a 10x10 maze. In both Maze2D and AntMaze, Gaussian noise is introduced at higher levels of $\sigma=0.5$, $1.0$, and $1.5$. Each environment is structured with distinct meta-train and meta-test tasks, ensuring that test tasks involve different goals from those seen during training. Further experimental details, including descriptions of other baselines, environment configurations (number of tasks, state representations, and reward setups), initial data collection processes, and hyperparameter settings, are provided in Appendix C.
@@ 5.3 Performance Comparison
We compare the proposed SISL with various baseline algorithms. Non-meta RL algorithms are trained directly on each test task for 0.5K iterations due to the absence of a meta-train phase. Meta-RL algorithms undergo meta-training for 10K iterations in Kitchen and Office, and 4K in Maze2D and AntMaze, followed by fine-tuning on test tasks for an additional 0.5K iterations. For SISL, the skill refinement interval $K_{\text{iter}}$ is set to 2K for Kitchen, Office, and AntMaze; 1K for Maze2D. To ensure a fair comparison, SISL counts each update process from its skill-improvement policy and high-level policy as one iteration. Table 1 presents the final average return across test tasks after the specified test iterations. The corresponding learning curves are provided in Appendix D.1 for a more detailed comparison. From the result, SAC and PEARL baselines, which do not utilize skills or offline datasets, perform poorly on long-horizon tasks, yielding a single result across all noise levels. In contrast, SPiRL, SiMPL, and SISL, which leverage skills, achieve better performance.
@@
SPiRL and SiMPL, however, show sharp performance declines as dataset noise increases. While both perform well with expert data, SiMPL struggles under noisy conditions due to instability in its task encoder $q_{e}$, sometimes performing worse than SPiRL. Here, the baseline results for Maze2D (Expert) are somewhat lower than those reported in the SiMPL paper. This discrepancy likely arises because, in constructing the offline dataset, we considered fewer tasks compared to SiMPL, resulting in trajectories that do not fully cover the map. Interestingly, minor noise occasionally boosts performance by introducing diverse trajectories that improve skill learning, a deteailed analysis of changes in state coverage is provided in Appendix D.1. In contrast, SISL demonstrates superior robustness across all evaluated environments, consistently outperforming baselines at varying noise levels. For example, in the Kitchen environment, SISL maintains strong performance under significant noise by effectively refining useful skills, while in Maze2D, higher noise levels lead to the improvement of diverse skills, achieving perfect task completion when $\sigma=1.5$. These results highlight SISL’s ability to discover improved behavior and refine robust skills, significantly enhancing meta-RL performance. Moreover, SISL excels with both noisy and expert data, achieving superior test performance by learning more effective skills.
Fig. 6 shows the learning progress during the meta-train/test phases for Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), highlighting the performance gap between SISL and other methods. The periodic drops in SISL’s high-level performance correspond to the reinitialization of $\pi_{h}$ every $K_{\text{iter}}$. Non-meta RL algorithms, including those with RND-based exploration, struggle with long-horizon tasks, while SPiRL and SiMPL show limited improvement due to their reliance on noisy offline datasets. In contrast, SISL’s self-improving skill refinement supports continuous skill improvement, resulting in superior meta-test performance. To further assess SISL’s robustness, we perform experiments with limited offline data, random noise injection, and diverse sub-optimal datasets in Appendix D, reflecting real-world challenges such as costly data collection and unstructured anomalies. Even under these conditions, SISL consistently outperforms the baselines, demonstrating strong robustness. Furthermore, we provide a computational cost analysis with SiMPL in Appendix E, and the result shows that SISL requires about 16% more computation time per iteration during meta-training, while the meta-test phase remains unchanged. Notably, even with extended training, SiMPL fails to achieve further performance gains, highlighting SISL’s advantage.
@@
Figure 7: Visualization of buffer mixing coefficient $\beta$ dynamics and refined skill evolution in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). In Kitchen, the refined skills at $t=15$ and $t=30$ during the microwave-opening task are depicted, while in Maze2D, trajectories using refined skills illustrate the process of progressively expanding to broader areas to solve tasks.
5.4 In-depth Analysis of the Proposed Skill Refinement Process
To analyze skill refinement and prioritization in more detail, Fig. 7 illustrates the evolution of the buffer mixing coefficient $\beta$ and skill refinement in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$). For Kitchen, the microwave-opening subtask is evaluated, while Maze2D focuses on navigation improvements. In the early stages (1K iterations for Kitchen, 0.5K for Maze2D), pretrained skills from the offline dataset are used without updates, resulting in poor performance, with the agent failing to grasp the handle in Kitchen and producing noisy trajectories in Maze2D. As training progresses, $\beta$ increases to shift contribution from offline data to newly collected high-quality data, providing task-relevant refinement. By design, the increase of $\beta$ tracks task-return improvement and serves as a soft curriculum that prevents abrupt distribution shift. At the same time, prioritized offline samples keep $\beta$ below 1, thereby ensuring generalization.
@@ 5.4 In-depth Analysis of the Proposed Skill Refinement Process
As a result, by iteration 5K in Kitchen, the agent learns to open the microwave, refining this skill to complete the task more efficiently by iteration 10K. In Maze2D, the agent explores more diverse trajectories over iterations, ultimately solving all training tasks by iteration 4K. These results highlight how SISL refines skills iteratively by leveraging prioritized data from offline and online buffers. As shown in Table 1, the learned skills generalize effectively to unseen test tasks, demonstrating SISL’s robustness and efficacy. While variations in return scales across tasks can introduce bias in reward model training, our experimental environments adopt a simple reward structure based on subtask completion, under which the reward model remains stable, as shown in Appendix F.4. For environments with more complex reward functions, per-task reward standardization can be considered. We provide $\beta$ trends and visualize refined skills, task-representation improvements, and policy skill composition in Appendix F. These analyses provide insights into SISL’s effectiveness in optimizing skill execution and enhancing task representation.
@@ 5.5 Ablation Studies
Figure 8: Component evaluation
Figure 9: Impact of prioritization temperature $T$
We evaluate the impact of SISL’s components and key hyperparameters in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), focusing on the effect of the prioritization temperature $T$. Additional analyses, including component evaluation and KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ for all noise levels, are in Appendix G.
Component Evaluation: To evaluate the importance of SISL’s components, we compare the meta-test performance of SISL with all components included against the following variations: (1) Without $\mathcal{B}_{\text{off}}$, relying solely on $\mathcal{B}_{\text{on}}^{i}$, to evaluate the influence of offline data in skill refinement; (2) Without $P_{\mathcal{B}_{\text{off}}}$, applying uniform sampling in $\mathcal{B}_{\text{off}}$ instead of maximum return relabeling, to evaluate the effect of prioritization; (3) Without $\mathcal{B}_{\text{on}}$, using only $\mathcal{B}_{\text{off}}$, to evaluate the contribution of high-quality samples obtained by $\pi_{\text{imp}}$ and $\pi_{h}$; and (4) Without $\pi_{\text{imp}}$, removing the skill-improvement policy, to evaluate whether exploration near the skill distribution discovers improved behaviors. Fig. 9 presents the comparison results, demonstrating significant performance drops when either buffer is removed, highlighting their critical role in effective skill discovery. Uniform sampling in $\mathcal{B}_{\text{off}}$ also reduces performance, underlining the importance of maximum return relabeling. Lastly, excluding $\pi_{\text{imp}}$ notably degrades performance, emphasizing the critical role of discovering improved behavior.
Prioritization Temperature $T$: The prioritization temperature $T$ adjusts the prioritization between online and offline buffers. Specifically, lower $T$ biases sampling toward high-return buffers, while higher $T$ results in uniform sampling. Fig. 9 illustrates the performance variations with different prioritization temperatures $T$. When $T=0.1$, performance degrades due to excessive focus on a single buffer, aligning with the trends observed in the component evaluation. Conversely, high $T=2.0$ also degrades performance by eliminating prioritization. These results highlight the importance of proper tuning: $T=1.0$ for Kitchen and $T=0.5$ for Maze2D achieve the best performance. Based on the result, we set $T$ approximately proportional to the high-return range of each environment, which consistently yielded the best performance while avoiding extensive tuning.
6 Limitation
Although SISL demonstrates strong performance, it has several limitations. First, although SISL achieves notable performance gains, its computation time per iteration increases by 16% over the baseline. This overhead mainly comes from training the skill model without freezing it during meta-training, but the performance table and ablation study confirm that this component is essential. (See Appendix E for details) Second, SISL requires fine-tuning during the meta-test phase for optimal performance, which introduces additional computational overhead. Addressing this through zero-shot skill adaptation could enhance its practicality, enabling transfer to new tasks without retraining. Future work in this direction could significantly improve SISL’s applicability in real-world scenarios.